# Bayesian Doubly Robust Causal Inference via Posterior Coupling [^1] October 16, 2025

### v2

@@ Abstract
Bayesian doubly robust (DR) causal inference faces a fundamental dilemma: joint modeling of outcome and propensity score suffers from the feedback problem where outcome information contaminates propensity score estimation, while two-step analysis sacrifices valid posterior distributions for computational convenience. We resolve this dilemma through posterior coupling via entropic tilting. Our framework constructs independent posteriors for propensity score and outcome models, then couples them using entropic tilting to enforce the DR moment condition. This yields the first fully Bayesian DR estimator with an explicit posterior distribution. Theoretically, we establish three key properties: (i) when the outcome model is correctly specified, the tilted posterior coincides with the original; (ii) under propensity score model correctness, the posterior mean remains consistent despite outcome model misspecification; (iii) convergence rates improve for nonparametric outcome models. Simulations demonstrate superior bias reduction and efficiency compared to existing methods. We illustrate practical advantages of the proposed method through two applications: sensitivity analysis for unmeasured confounding in antihypertensive treatment effects on dementia, and high-dimensional confounder selection combining shrinkage priors with modified moment conditions for right heart catheterization mortality. We provide an R package implementing the proposed method.
Keywords: Bayesian inference, Double robustness, Efficiency improvement, Entropic tilting, Propensity score
@@ Introduction
In previous studies, Bayesian DR estimation methods have been proposed while avoiding the cutting feedback problem. Saarela et al. (2016) proposed a Bayesian DR estimator by introducing a weighting loss function for the estimator, combined with the Bayesian bootstrap. This approach can be interpreted as defining a loss function for a pseudo-population in which confounding effects are removed (Hernán and Robins, 2020). While this method represents an important contribution from the Bayesian perspective, it does not yield an explicit posterior distribution. Antonelli et al. (2022); Luo et al. (2023) also proposed another Bayesian DR method that incorporate propensity score information in a thoughtful way; however, this approach similarly lack tractable expressions for the posterior distribution.
In this manuscript, we propose a novel framework for Bayesian DR inference that provides an explicit posterior distribution. Our key innovation is to couple separate (independent) posterior distributions for the outcome (i.e., excluding propensity score information) and propensity score models. In our proposed method, propensity score information is incorporated into the outcome model using entropic tilting (ET) (Jaynes, 1957; Tallman and West, 2022) based on a moment condition that corresponds to the augmentation term in the DR estimator (Bang and Robins, 2005), ensuring double robustness. While Yiu et al. (2020) employ a similar tilting idea, our approach differs in yielding an explicit posterior distribution with a notably simple moment condition. The recent work by Breunig et al. (2025) also develops a nonparametric Bayesian DR method, but requires auxiliary data.
@@
To address this, Saarela et al. (2016) and Breunig et al. (2025) consider modeling the joint distribution for the outcome and propensity score models, ensuring that the outcome model does not include information from the propensity score model. We follow this idea in the construction of the general posterior introduced in Section 3. The critical difference between our proposed method and previous works on Bayesian doubly robust estimation (Saarela et al., 2016; Yiu et al., 2020; Antonelli et al., 2022; Luo et al., 2023; Breunig et al., 2025) is that our approach constitutes a fully Bayesian estimation. Specifically, our method involves constructing the (general) posterior distribution of $\bm{\alpha}$ and $\bm{\beta}$, separately, and then modifying it through constraint to ensure doubly robustness.
@@
$\ell(\bm{\beta})$ $=\exp\left\{-nf_{n}(\bm{\alpha})\right\},\ \ \ \ f_{n}(\bm{\alpha})=\frac{1}{n}\sum_{i=1}^{n}f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta}),$ (3.1)
$\ell(\bm{\alpha})$ $=\exp\left\{-nf_{n}(\bm{\beta})\right\},\ \ \ \ f_{n}(\bm{\beta})=\frac{1}{n}\sum_{i=1}^{n}f(A_{i}\mid\bm{X}_{i};\bm{\alpha})$ (3.2)
where $\bm{\alpha}\in\Theta_{\bm{\alpha}}$ and $\bm{\beta}\in\Theta_{\bm{\beta}}$. Note that when $nf_{n}(\bm{\alpha})$ and $nf_{n}(\bm{\beta})$ are negative log-likelihood of a unit sample, both (3.1) and (3.2) reduce to the standard likelihood functions. The outcome model (3.1) could be constructed “model-free” squared loss function, $f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta})=\omega\left(Y_{i}-m_{A_{i}}(\bm{X}_{i};\bm{\beta})\right)^{2}$, where $\omega$ is a learning rate (e.g. Bissiri et al., 2016; Wu and Martin, 2023). For the propensity score model (3.2), a standard logistic regression model for the propensity score is equivalent to specifying $f(A_{i}\mid\bm{X}_{i};\bm{\alpha})=A_{i}\log\{e(\bm{X}_{i};\bm{\alpha})\}+(1-A_{i})\log\{1-e(\bm{X}_{i};\bm{\alpha})\}$ with propensity score $e(\bm{X}_{i};\bm{\alpha})$. It can also be derived from the covariate balancing propensity score conditions (Imai and Ratkovic, 2014; Orihara et al., 2024). In the following discussions, we assume that global maximizers of (3.1) and (3.2) exist, denoted by $\bm{\alpha}^{*}$ and $\bm{\beta}^{*}$, respectively.
@@
To construct doubly robust posterior, we couple information of two posteriors based on outcome and propensity score models. Specifically, we employ the entropic tilting (Jaynes, 1957; Tallman and West, 2022), to obtain constraint posterior distribution under pre-specified moment conditions. In particular, we employ the constraint $B_{n}(\bm{\alpha},\bm{\beta})=0$, where
@@
In our proposed method, the concept is similar to Saarela et al. (2016) and Luo et al. (2023), which tailor exponential tilting methods. In both methods, moment conditions such as (3.4) play an important role in constructing a doubly robust estimator. Their approaches are primarily based on non-Bayesian perspectives, with efforts to harmonize them with Bayesian contexts. In contrast, our proposed method relies solely on Bayesian perspectives. This is one of the important differences between these methods, and it makes our proposed method attractive from a practical perspective when applying existing Bayesian methods.
@@
The above method is based on the kernel smoothing updating of particles (Liu and West, 2001). A notable feature of the above algorithm requires generating random samples from the original posterior as initial particles. Also, it does not require evaluation of the value of original posterior, but it only needs the evaluation of the doubly robust constraint $B_{n}(\bm{\alpha},\bm{\beta})$. When Algorithm 2 is terminated at $t=t_{0}$, $\lambda_{(t_{0})}$ will be a desirable tilting parameter and $(\bm{\alpha}_{t}^{(s)},\bm{\beta}_{t}^{(s)})$ with equal weights are samples from the tilted posterior.
@@
We here discuss the theoretical properties of the tilted posterior and the doubly robustness of the posterior mean (3.8) under the tilted posterior. First, we show the behavior of the tilted posterior when the outcome model is correctly specified, as given in the following lemma.
@@
aAssuming regularity conditions in Appendix A. When the outcome model is correctly specified, it holds the following property under $n\to\infty$:
@@
Furthermore, the posterior mean (3.8) has a doubly robustness property as shown in the following theorem:
@@
aAssuming regularity conditions in Appendix A. When either the outcome model or the propensity score model is correctly specified, then it holds under $n\to\infty$ that
@@
In Saarela et al. (2016), the authors proposed a Bayesian bootstrap-based estimator. Let $\xi=(\xi_{1},\dots,\xi_{n})$ denote the weights for the Bayesian bootstrap, with $\Xi\mid D\sim\mathrm{Dir}(1,\dots,1)$. Using the sampled weights $\xi$, a doubly robust estimator weighted by $\xi$ is constructed:
@@
In Yiu et al. (2020) and Luo et al. (2023), a novel Bayesian method was proposed by combining empirical likelihood with the Bayesian bootstrap, where the ATE $\tau$ is the target estimand. In their framework, the moment condition (3.6), or analogous moment conditions, are essential from a frequentist standpoint:
$\sum_{i=1}^{n}\exp\left\{\hat{\lambda}\left(\frac{A_{i}-e(\bm{X}_{i};\bm{\alpha})}{e(\bm{X}_{i};\bm{\alpha})(1-e(\bm{X}_{i};\bm{\alpha}))}\left(Y_{i}-m_{A_{i}}(\bm{X}_{i};\bm{\beta})\right)\right)\right\}$
$\hskip 14.22636pt\times\frac{A_{i}-e(\bm{X}_{i};\bm{\alpha})}{e(\bm{X}_{i};\bm{\alpha})(1-e(\bm{X}_{i};\bm{\alpha}))}\left(Y_{i}-m_{A_{i}}(\bm{X}_{i};\bm{\beta})\right)=0,$
where $\lambda$ denotes a Lagrange multiplier, and $\hat{\lambda}$ denotes its estimator. The origin of $\lambda$ differs from that in our proposed method; however, in both their methods and our proposed approach, the estimation (or adjustment) of $\lambda$ plays an essential role in achieving double robustness.
@@
To confirm performance of our proposed procedure, we conducted simulation experiments. The iteration time of all simulations was 2000. A simulation experiment under a high-dimensional setting is presented in Appendix D.
@@
We compared four methods: one non-Bayesian DR estimator proposed by Bang and Robins (2005), one Bayesian G-formula based method (Daniels et al., 2023), two Bayesian DR estimator proposed by Saarela et al. (2016) and Luo et al. (2023), and the proposed Bayesian DR estimator using ET.
@@ 5.3.1 Remaining bias modification
To address this problem, we propose the “sample pruning” algorithm. When updating $\lambda$ in the SMC algorithm, we discard samples with small sampling weights. As a result, the remaining samples are more concentrated around $B_{n}\approx 0$ without extreme sampling weights. The results of the sample pruning algorithm are presented in the last row of Table 1. The remaining bias is clearly diminished.
@@
As mentioned in Section 3.5, the proposed method can improve the convergence rate of nonparametric methods. The results are summarized in Table 2. As expected, the proposed method improves upon the results of the G-formula, especially in small-sample situations. This is because our proposed method balances the moment condition (3.4) using ET, and this term diminishes (goes to 0) asymptotically. Additionally, the simulation results suggest that using ET improves efficiency even when the propensity score model is misspecified.
These results suggest that detecting a valid outcome model is an initially important task, and identifying a valid propensity score model may help improve the efficiency of the ATE estimation.
@@
We use the HRS data available from the Health and Retirement Study website (https://hrsdata.isr.umich.edu/data-products/). For our analysis, data from 2010 to 2016 were used at two-year intervals. Specifically, we defined the baseline characteristics based on the first two, and the exposure and outcome variables based on the latter two. Subjects who had dementia, Alzheimer’s disease, memory problems, or who received any antihypertensive treatment at baseline were excluded from the analysis. Additionally, subjects aged 15 to 65 years with systolic blood pressure greater than 130 at baseline were included in the analysis. The exposure variable was defined as “using any antihypertensive treatment and having hypertension” or not. Under this definition, we can evaluate the effect of antihypertensive treatment for hypertension. The outcome variable was defined as “dementia, Alzheimer’s disease, or memory problems” or not. Confounders were selected based on Bennett et al. (2025): baseline SBP, age, sex, education, race, body mass index, apolipoprotein E $\varepsilon$4 allele status, smoking, depression, coronary heart disease, stroke, and diabetes.
@@
${\rm E}\left[{\rm E}\left[Y\mid A=1,\bm{X}\right]-{\rm E}\left[Y\mid A=0,\bm{X}\right]\right]=\tau-\xi,$
@@
and $\Delta_{a}(\bm{X})={\rm E}\left[Y_{a}\mid A=1,\bm{X}\right]-{\rm E}\left[Y_{a}\mid A=0,\bm{X}\right]$. Note that the assumption of strongly ignorable treatment assignment corresponds to $\xi=0$. A simple but important case is when $\Delta_{a}(\bm{X})\equiv\Delta$, which corresponds to the situation where $\bm{X}$ provides no information about unmeasured confounders (Daniels et al., 2023). In this case, $\xi=\Delta$, and $\xi$ serves as the sole sensitivity parameter.
To consider the sensitivity parameter, we adopt an outcome model, $m_{A}(\bm{X};\bm{\beta},\xi)=m_{A}(\bm{X};\bm{\beta})-A\xi$. Let $f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta},\xi)$ be the corresponding negative log-likelihood function of the outcome model including $\xi$. Then, the joint posterior of $(\bm{\alpha},\bm{\beta})$ is proportional to
$\pi(\bm{\alpha})\pi(\bm{\beta})\prod_{i=1}^{n}\exp\left\{-f(A_{i}\mid\bm{X}_{i};\bm{\alpha})\right\}\int\exp\left\{-f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta},\xi)\right\}g(\xi)d\xi,$ (6.1)
where $g(\xi)$ is a density function of $\xi$. Then, by using the posterior samples from the original posterior $p_{n}(\bm{\alpha},\bm{\beta}\mid D)$ with the base model $m_{A}(\bm{X};\bm{\beta})$, the posterior samples from the above posterior can be generated by an importance sampling with important weight proportional to
$\prod_{i=1}^{n}\frac{\int\exp\left\{-f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta},\xi)\right\}g(\xi)d\xi}{\exp\left\{-f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta})\right\}}.$
@@
$w^{(s)}\equiv\prod_{i=1}^{n}\frac{M^{-1}\sum_{m=1}^{M}\exp\left\{-f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta}^{(s)},\xi^{(m)})\right\}}{\exp\left\{-f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta}^{(s)})\right\}}.$
@@
Given the posterior samples of $(\bm{\alpha},\bm{\beta})$ under unmeasured confounders, we can also obtain the doubly robust posterior samples of the ATE through posterior coupling, described in Algorithms 1 or 2.

### v3

# Bayesian Doubly Robust Causal Inference via Posterior Coupling [^1] January 2, 2026

@@ Abstract
Bayesian doubly robust (DR) causal inference faces a fundamental dilemma: joint modeling of outcome and propensity score suffers from the feedback problem where outcome information contaminates propensity score estimation, while two-step inference sacrifices valid posterior distributions for computational convenience. We resolve this dilemma through posterior coupling via entropic tilting. Our framework constructs independent posteriors for propensity score and outcome models, then couples them using entropic tilting to enforce the DR moment condition. This yields the first fully Bayesian DR estimator with an explicit posterior distribution. Theoretically, we establish three key properties: (i) when the outcome model is correctly specified, the tilted posterior coincides with the original; (ii) under propensity score model correctness, the posterior mean remains consistent despite outcome model misspecification; (iii) convergence rates improve for nonparametric outcome models. Simulations demonstrate superior bias reduction and efficiency compared to existing methods. We illustrate practical advantages of the proposed method through two applications: sensitivity analysis for unmeasured confounding in antihypertensive treatment effects on dementia, and high-dimensional confounder selection combining shrinkage priors with modified moment conditions for right heart catheterization mortality. We provide an R package implementing the proposed method.
Keywords: Bayesian inference, Efficiency improvement, Entropic tilting, Propensity score, Sequential Monte Carlo
@@ Introduction
In previous studies, Bayesian DR estimation methods have been proposed while avoiding the cutting feedback problem. Saarela et al. (2016) proposed a Bayesian DR estimator by introducing a weighting loss function for the estimator, combined with the Bayesian bootstrap. This approach can be interpreted as defining a loss function for a pseudo-population in which confounding effects are removed (Hernán and Robins, 2020). While this method represents an important contribution from the Bayesian perspective, it does not yield an explicit posterior distribution. Antonelli et al. (2022) also proposed another Bayesian DR method that incorporate propensity score information in a thoughtful way; however, this approach similarly lack tractable expressions for the posterior distribution.
In this manuscript, we propose a novel framework for Bayesian DR inference that provides an explicit posterior distribution. Our key innovation is to couple separate (independent) posterior distributions for the outcome (i.e., excluding propensity score information) and propensity score models. In our proposed method, propensity score information is incorporated into the outcome model using entropic tilting (ET) (Jaynes, 1957; Tallman and West, 2022) based on a moment condition that corresponds to the augmentation term in the DR estimator (Bang and Robins, 2005), ensuring double robustness. While Yiu et al. (2020) and Luo et al. (2023) employ a similar tilting idea, our approach differs in that it yields an explicit posterior distribution with a relatively simple moment condition and procedure. The recent work by Breunig et al. (2025) also develops a nonparametric Bayesian DR method, but requires auxiliary data.
@@
To address this, Saarela et al. (2016) considers modeling the joint distribution for the outcome and propensity score models, ensuring that the outcome model does not include information from the propensity score model. We follow this idea in the construction of the general posterior introduced in Section 3. The critical difference between our proposed method and previous works on Bayesian DR estimation (Saarela et al., 2016; Yiu et al., 2020; Antonelli et al., 2022; Luo et al., 2023; Breunig et al., 2025) is that our approach constitutes a fully Bayesian estimation. Specifically, our method involves constructing the (general) posterior distribution of $\bm{\alpha}$ and $\bm{\beta}$, separately, and then modifying it through constraint to ensure double robustness.
@@
$\ell(\bm{\beta})$ $=\exp\left\{-nf_{n}(\bm{\beta})\right\},\ \ \ \ f_{n}(\bm{\beta})=\frac{1}{n}\sum_{i=1}^{n}f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta}),$ (3.1)
$\ell(\bm{\alpha})$ $=\exp\left\{-nf_{n}(\bm{\alpha})\right\},\ \ \ \ f_{n}(\bm{\alpha})=\frac{1}{n}\sum_{i=1}^{n}f(A_{i}\mid\bm{X}_{i};\bm{\alpha})$ (3.2)
where $\bm{\alpha}\in\Theta_{\bm{\alpha}}$ and $\bm{\beta}\in\Theta_{\bm{\beta}}$. Note that when $nf_{n}(\bm{\alpha})$ and $nf_{n}(\bm{\beta})$ are negative log-likelihood of a unit sample, both (3.1) and (3.2) reduce to the standard likelihood functions. The outcome model (3.1) could be constructed “model-free” squared loss function, $f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta})=\omega\left(Y_{i}-m_{A_{i}}(\bm{X}_{i};\bm{\beta})\right)^{2}$, where $\omega$ is a learning rate (e.g. Bissiri et al., 2016; Wu and Martin, 2023). For the propensity score model (3.2), a standard logistic regression model for the propensity score is equivalent to specifying $f(A_{i}\mid\bm{X}_{i};\bm{\alpha})=A_{i}\log\{e(\bm{X}_{i};\bm{\alpha})\}+(1-A_{i})\log\{1-e(\bm{X}_{i};\bm{\alpha})\}$ with propensity score $e(\bm{X}_{i};\bm{\alpha})$. It can also be derived from the covariate balancing propensity score conditions (Imai and Ratkovic, 2014; Orihara et al., 2025). In the following discussions, we assume that global maximizers of (3.1) and (3.2) exist, denoted by $\bm{\alpha}^{*}$ and $\bm{\beta}^{*}$, respectively.
@@
To construct DR posterior, we couple information of two posteriors based on outcome and propensity score models. Specifically, we employ the entropic tilting (Jaynes, 1957; Tallman and West, 2022), to obtain constraint posterior distribution under pre-specified moment conditions. In particular, we employ the constraint $B_{n}(\bm{\alpha},\bm{\beta})=0$, where
@@
The above method is based on the kernel smoothing updating of particles (Liu and West, 2001). A notable feature of the above algorithm requires generating random samples from the original posterior as initial particles. Also, it does not require evaluation of the value of original posterior, but it only needs the evaluation of the DR constraint $B_{n}(\bm{\alpha},\bm{\beta})$. When Algorithm 2 is terminated at $t=t_{0}$, $\lambda_{(t_{0})}$ will be a desirable tilting parameter and $(\bm{\alpha}_{t}^{(s)},\bm{\beta}_{t}^{(s)})$ with equal weights are samples from the tilted posterior.
@@
We here discuss the theoretical properties of the tilted posterior and the double robustness of the posterior mean (3.8) under the tilted posterior. First, we show the behavior of the tilted posterior when the outcome model is correctly specified, as given in the following lemma.
@@
Assuming regularity conditions in Appendix A. When the outcome model (3.1) is correctly specified, it holds the following property under $n\to\infty$:
@@
Furthermore, the posterior mean (3.8) has a double robustness property as shown in the following theorem:
@@
Assuming regularity conditions in Appendix A. When either the outcome model (3.1) or the propensity score model (3.2) is correctly specified, then it holds under $n\to\infty$ that
@@
Remark 1.
For our proposed methodology, model correctness is related to the assumed (pseudo-) likelihood. For instance, when we assume the full likelihood for the outcome model, a fully correct specification of the likelihood function is necessary. Whereas, when we assume only the mean construction for the outcome model based on general Bayes, correctness of the outcome mean construction is sufficient. This point is different from traditional double robustness (Tsiatis, 2006) and previous Bayesian DR methods (Saarela et al., 2016; Luo et al., 2023).
Remark 2.
When both models are correctly specified, the posterior distribution after posterior coupling is the same as the posterior before coupling (from Lemma 1). Therefore, the efficiency is derived from the outcome model, making it more efficient than traditional DR estimators. This point is a benefit of our proposed methodology and can be confirmed in the next simulation experiments.
@@
In our proposed method, the concept is similar to Saarela et al. (2016) and Luo et al. (2023). In both methods, moment conditions such as (3.4) play an important role in constructing a DR estimator. Their approaches are primarily based on non-Bayesian perspectives, with efforts to harmonize them with Bayesian contexts.
In Saarela et al. (2016), the authors proposed a Bayesian bootstrap-based estimator. Let $\xi=(\xi_{1},\dots,\xi_{n})$ denote the weights for the Bayesian bootstrap, with $\Xi\mid D\sim\mathrm{Dir}(1,\dots,1)$. Using the sampled weights $\xi$, a DR estimator weighted by $\xi$ is constructed:
@@
In Yiu et al. (2020) and Luo et al. (2023), a novel Bayesian method was proposed by tailoring the empirical likelihood framework (Owen, 2001), where the ATE $\tau$ is the target estimand. Here, to clarify the discussion, we consider the true outcome model as
$Y_{i}=\tau A_{i}+h_{0}(\bm{X}_{i};\bm{\beta}^{0})+\varepsilon_{i},$
where $\varepsilon_{i}$ is an i.i.d. error. To estimate the parameters, we consider the working model
[…]
The estimated propensity score is plugged into the model; however, simultaneous estimation can also be considered (Yiu et al., 2020). To estimate the parameters, we consider the following estimating equation:
[…]
By subtracting the 3rd row of (4.5) from the 1st row, we obtain
$\widehat{\tilde{\tau}}=\frac{\sum_{i=1}^{n}(A_{i}-e_{i}(\hat{\bm{\alpha}}))(Y_{i}-h_{1}(\bm{X}_{i};\hat{\bm{\beta}})-\hat{\phi}e_{i}(\hat{\bm{\alpha}}))}{\sum_{i=1}^{n}A_{i}(A_{i}-e_{i}(\hat{\bm{\alpha}}))}$,
and this estimator is related to the DR estimator (2.2). This is because, from the true outcome model,
$\widehat{\tilde{\tau}}\approx\tau+\frac{\sum_{i=1}^{n}(A_{i}-e_{i}(\hat{\bm{\alpha}}))(h_{0}(\bm{X}_{i};\bm{\beta}^{0})-h_{1}(\bm{X}_{i};\hat{\bm{\beta}})-\hat{\phi}e_{i}(\hat{\bm{\alpha}}))}{\sum_{i=1}^{n}A_{i}(A_{i}-e_{i}(\hat{\bm{\alpha}}))}$.
The estimator has a form similar to (2.2), and the second term clearly converges to zero asymptotically when either the propensity score or a part of the outcome model (i.e., $h_{1}(\bm{X}_{i};\bm{\beta})$) is correctly specified. Using (4.5), we can consider the empirical likelihood estimator. Specifically, the estimated empirical likelihood function can be expressed as
$p_{i}(\bm{\beta},\phi,\tilde{\tau})=\frac{\exp\left\{\hat{\lambda}^{\top}U_{i}(\bm{\beta},\phi,\tilde{\tau})\right\}}{\sum_{j=1}^{n}\exp\left\{\hat{\lambda}^{\top}U_{j}(\bm{\beta},\phi,\tilde{\tau})\right\}}$,
where $\lambda$ is a Lagrange multiplier. Using $p_{i}$, we can consider the posterior distribution
[…]
Based on the posterior distribution, one can obtain samples that possess the double robustness property. The origin of $\lambda$ differs from that in our proposed method; however, in both their methods and our proposed approach, the estimation (or adjustment) of $\lambda$ plays an essential role in achieving double robustness. Compared with this procedure, we consider our proposed methodology more straightforward, as it simply combines the posterior distributions for the propensity score and outcome models.
@@
To confirm performance of our proposed procedure, we conducted simulation experiments. The iteration time of all simulations was 2000. Some additional simulation experiments are presented in Appendix D.
@@
We compared four methods: one non-Bayesian DR estimator proposed by Bang and Robins (2005), one Bayesian G-formula based method (Daniels et al., 2023), two Bayesian DR estimator proposed by Saarela et al. (2016) and Luo et al. (2023), and the proposed Bayesian DR estimator using posterior coupling.
@@ 5.3.1 Remaining bias modification
To address this problem, we propose the “sample pruning” algorithm. When updating $\lambda$ in the SMC algorithm, we discard samples with small sampling weights (i.e., $\exp\left\{\lambda B_{n}\right\}$). As a result, the remaining samples are more concentrated around $B_{n}\approx 0$ without extreme sampling weights. The results of the sample pruning algorithm are presented in the last row of Table 1. The remaining bias is clearly diminished.
@@
As mentioned in Section 4.2, the proposed method can improve the convergence rate of nonparametric methods. The results are summarized in Table 2. As expected, the proposed method improves upon the bias and variance of the results of the G-formula, especially in small-sample situations. This is because our proposed method balances the moment condition (3.4) using ET, and this term diminishes (goes to 0) asymptotically. Additionally, the simulation results suggest that using ET improves efficiency even when the propensity score model is misspecified.
These results suggest that detecting a valid outcome model is an initially important task, and incorporating a propensity score model (information) may help improve the efficiency of the ATE estimation.
@@
We use the HRS data available from the Health and Retirement Study website (https://hrsdata.isr.umich.edu/data-products/). For our analysis, data from 2010 to 2016 were used at two-year intervals. Specifically, we defined exposure and baseline characteristics using the 2010 and 2012 surveys, and assessed outcome variables using the data collected in the survey conducted four years later. Subjects who had dementia, Alzheimer’s disease, memory problems, or who received any antihypertensive treatment at baseline were excluded from the analysis. Additionally, subjects aged 65 years or older with systolic blood pressure greater than 130 at baseline were included in the analysis. The exposure variable was defined as “using any antihypertensive treatment and having hypertension” or not. The outcome variable was defined as “dementia, Alzheimer’s disease, or memory problems” or not. Confounders were selected based on Bennett et al. (2025): baseline SBP, age, sex, education, race, body mass index, apolipoprotein E $\varepsilon$4 allele status, smoking, depression, coronary heart disease, stroke, and diabetes. Under these definitions, we can evaluate the effect of antihypertensive use on dementia incidence.
@@
${\rm E}\left[{\rm E}\left[Y\mid A=1,\bm{X}\right]-{\rm E}\left[Y\mid A=0,\bm{X}\right]\right]=\tau+\xi,$
@@
and $\Delta_{a}(\bm{X})={\rm E}\left[Y_{a}\mid A=1,\bm{X}\right]-{\rm E}\left[Y_{a}\mid A=0,\bm{X}\right]$ (more details, Hu et al. 2022). Note that the assumption of strongly ignorable treatment assignment corresponds to $\xi=0$. A simple but important case is when $\Delta_{a}(\bm{X})\equiv\Delta$, which corresponds to the situation where $\bm{X}$ provides no information about unmeasured confounders (Daniels et al., 2023). In this case, $\xi=\Delta$, and $\xi$ serves as the sole sensitivity parameter.
To consider the sensitivity parameter, we adopt an outcome model, $m_{A}(\bm{X};\bm{\beta},\xi)=m_{A}(\bm{X};\bm{\beta})+A\xi$. Let $f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta},\xi_{i})$ be the corresponding negative log-likelihood function of the outcome model including $\xi_{i}$. Then, the joint posterior of $(\bm{\alpha},\bm{\beta})$ is proportional to
$\pi(\bm{\alpha})\pi(\bm{\beta})\prod_{i=1}^{n}\exp\left\{-f(A_{i}\mid\bm{X}_{i};\bm{\alpha})\right\}\int\exp\left\{-f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta},\xi_{i})\right\}g(\xi_{i})d\xi_{i},$ (6.1)
where $g(\xi_{i})$ is a density function of $\xi_{i}$. Then, by using the posterior samples from the original posterior $p_{n}(\bm{\alpha},\bm{\beta}\mid D)$ with the base model $m_{A}(\bm{X};\bm{\beta})$, the posterior samples from the above posterior can be generated by an importance sampling with important weight proportional to
$\prod_{i=1}^{n}\frac{\int\exp\left\{-f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta},\xi_{i})\right\}g(\xi_{i})d\xi_{i}}{\exp\left\{-f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta})\right\}}.$
@@
(Importance sampling approach for sensitivity analysis) For $s=1,\ldots,S$, generate $M$ samples from $g(\cdot)$, denoted by $\xi^{(m)}_{i}$ for $m=1,\ldots,M$ and compute the weight as
$w^{(s)}\equiv\prod_{i=1}^{n}\frac{M^{-1}\sum_{m=1}^{M}\exp\left\{-f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta}^{(s)},\xi^{(m)}_{i})\right\}}{\exp\left\{-f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta}^{(s)})\right\}}.$
@@
Note that in the above sensitivity analysis case ($\xi=\Delta$), $\xi_{i}$ can be considered as $\xi_{i}\equiv\xi$. Another importance weighting derivation strategy is discussed in Appendix E.
Given the posterior samples of $(\bm{\alpha},\bm{\beta})$ under unmeasured confounders, we can also obtain the DR posterior samples of the ATE through posterior coupling, described in Algorithms 1 or 2.